silver metal crystalline with a face centred cubic lattice the lenght of unit cell in  found tobethe value 4.077x10-8m.ca;culate the atomic radius and density of silver {atomic mass of silver(ag)=108units,NA=6.022x1023atoms /mol}

Grade:12

2 Answers

Sunil Kumar FP
9 years ago
we have
density of the fcc=z*M/a^3*Na
z=no of atom fro fcc it is 4
M= molecular mass
Na=avagadro number
a=edge length
for face centered cubic unit cell=sqrt2 *a=4r,r=1.414*4.077/4=1.44*10^-8m
we have density=4*108/6.022*10^23*(4.077*10^-8)^3
=1.05*10^5g/cm^3
Sunil Kumar FP
9 years ago
“silver metal crystalline with a face centred cubic lattice the lenght of unit cell in found tobethe value 4.077x10-8m.ca;culate the atomic radius and density of silver {atomic mass of silver(ag)=108units,NA=6.022x1023atoms /mol}“

I think in the length of the unit cell you should have edge length=4.077A0